O.K. what is (9, -1) and (6 , -2) in slope intercept form?
Answer by scott8148(6628)

the slope of the line through the points is [-2-(-1)]/[6-9] or 1/3

using point-slope __ y-(-1)=(1/3)(x-9) __ y+1=(1/3)x-3 __ y=(1/3)x-4